OpenStudy (anonymous):

Simplify the following expression. (3mn^4)^3 / (3m^3n)^4 A. 3/m^9n^8 B. n8/3m^9 C. 1/3m^9n^8 D. 3m^9n^8

OpenStudy (anonymous):

\[\frac{ 3^{3}m ^{3}n ^{4} }{ 3^{4}m ^{12}n ^{4} }=\frac{ 1 }{ 3m ^{9} }\]

OpenStudy (anonymous):

Which option would it be?

OpenStudy (anonymous):

sorry didnt see n^4(3) \[\frac{ n ^{8} }{ 3m ^{9} }\]=final answer

OpenStudy (anonymous):

Thanks!! So C right?

OpenStudy (anonymous):

not C. look at my final answer again

OpenStudy (anonymous):

Okay! Thanks, I didn't scroll down and see your final answer woops. But now I've got it